MODS1 Commissioning Run 7: UTC 2011 April 15

Observer(s): R. Pogge, R. Stoll (OSU), E. Skillman (UMinn), M. Fumana (INAF), M. Wagner, O. Kuhn (OSU/LBTO)
Telescope Operator: G. Bechetti (LBTO)
AO Support: Remote from Tucson as needed
Telescope Support: J. Urban (LBTO)
Instrument Support: D. Officer (LBTO)

Summary

Lighter winds and wamer than last night, but with some cirrus during the first and last parts of the night, clear after midnight for a couple of hours. We worked all night, but seeing was highly variable, between 0.7 and 3 arcsec at times. We did what we could to exercise the system. Still problems with the GCS capturing background regions during bright moonlight when seeing or clouds cause the guide star to fade. This caused us lost observations and great frustration.

Got successful tests of the acquisition and observation scripts, especially got excellent data and mileage on target acquisition onto slits.

Lost about 30min to a primary mirror panic.

Preparations

LN2 fill as normal. Minor software tweaks to scripting based on previous nights' testing.

Doug Miller has the new Shack-Hartmann reconstructor program ready for testing on-sky tonight. We plan to check it out first thing, and if stable we'll try to use it all night, or revert if there are issues. On subsequent nights a full set of WFS measurement scripts will be run from downtown to finalize (we hope) the remaining loose-end of the AGw system characterization.

Details

All times are UTC unless otherwise noted.

2300 - Wind has calmed considerably since yesterday, some thin banks of cirrus drifting in from the NW during the late afternoon. Lots of haze and dust in the surrounding valleys leftover from the high winds across the state last night. Less hazy in the afternoon than it was this morning at sunrise, but still a little hazy.

0150 - Open and equilibrating. Ventilation running, mirrors are at +2°C, outside is 0.2°C, chamber air temp is falling quickly, so we have a little ways to go. Bands of cirrus coming down from the NNW; the IR satellite animation shows they could be with us for the first part of the night.

Pointing set
  IE=-84
  CA=+90
Set collimation n BS9139

Setup to run on the standard star PG0823+546 to test a prototype acqslit script (align target to long slit). worked well after the inevitable oopses.

0345 - Cirrus blowing over, ready to start reconstructor tests...

Doug Miller dialed in to setup the new reconstructor test. On-axis it worked fine relative to the old reconstructor, but with a different zero-point.

Off-axis star 4.5-arcmin in Y from the star resulted in some problems with the new reconstructor diverging. Doug gathered data to try to diagnose the situation.

Had to reboot the AzCamServer for the AGw guide camera at one point - this is the same as the problem described in IssueTrak #3252. The problem is still with us (no sign much has been done to address it).

Seeing now awful, 2-3arcsec, so end of what can be done with WFS check out for now.

After a while the seeing started to improved down to 1 to 1.5-arcsec. Ran an acquisition at a high PA of an HII region field in NGC 5194. This exercised the acquisition scripts, and was a setup in a challenging crowded field. Took a 300s dual-grating shot, then 3x600s, then a 600s in red-only grating mode. Very nice data, excellent tracking and stability through the sequence.

0750 - Did an MOS acquisition sequence on Abell 1689 to test the revised modsalign.py script from the previous run. Time from end of the preset to opening the shutter on the spectra with the mask aligned was 13 minutes. Took thru-slit images and field images to give to modsalign.py for processing, executed the offset, took a confirmatory through-slit image to verify we were centered (we were, very well), then fired an observing script that configured the instrument for dual-beam grating spectroscopy and fired off 3x600s integrations. The results were quite beautiful, and a great test of the mask alignment code. Next step will be to make it more user friendly and write it up.

0930 - While trying the challenging faint-target slit acquisition on an r=22 faint galaxy in bright moon. We were able to see the target, but in trying to setup, the GCS kept failing by grabbing the bright bar in the guider subfield when the seeing puffed up (seeing was 1.5-2 arcsec during the acquisition). This is the silly "brightness veto" problem we've been reporting. We find the situation intolerable in bright moonlight. It made it impossible to observe this target, requiring us to abort the exposures (i.e., stop and readout what we had) when it sent the guide star off into the weeds.

1002 - While doing a preset the left primary mirror panicked. Took Geno 20 minutes to recover

Once the moon was down, we shot a faint (r=22.7) galaxy, a very challenging acquisition. We were essentially running nearly blind in puffy seeing, but in 600s we got clear detections of emission lines but no continuum. A good demonstration of the robustness of our procedures.

Typical long-slit acquisition times are being analyzed from the runtime logs.

Shot a few more standard stars, working on some multi-part observing scripts.

1215 - Twilight coming up, sending it to horizon to close up.

-- RichardPogge - 14 Apr 2011
Topic revision: r4 - 15 Apr 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback